OpenStudy (anonymous):

Square root 4x+1 - square root x-2=3

OpenStudy (anonymous):

\[\sqrt{4x+1}-\sqrt{x-2}=3\] \[\sqrt{41+1}=3+\sqrt{x-2}\] now you are going to have to square twice

OpenStudy (anonymous):

the first time you get \[4x+1=9+x-2+6\sqrt{x-2}\]

OpenStudy (anonymous):

or \[4x+1=7+x+6\sqrt{x-2}\] now isolate the radical

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

okay lets go slow

OpenStudy (anonymous):

start here \[\sqrt{4x+1}=3+\sqrt{x-2}\] and square both sides to get rid of the radical we get \[(\sqrt{4x+1})^2=\left(3+\sqrt{x-2}\right)^2\]

OpenStudy (anonymous):

the left hand side is \(4x+1\) because you are squaring the square root

OpenStudy (anonymous):

can't you leave the √x-2 on the left side?

OpenStudy (anonymous):

you can, but it will make a big big mess if you do

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

once we have \[(\sqrt{4x+1})^2=\left(3+\sqrt{x-2}\right)^2\] we get \[4x+1=(3+\sqrt{x-2})(3+(\sqrt{x-2})\] and you have to multiply out

OpenStudy (anonymous):

multiply out on the right using the distributive law gives \[4x+1=3\times 3+3\times \sqrt{x-2}+3\times \sqrt{x-2}+\sqrt{x-2}\times \sqrt{x-2}\] or \[4x+1=9+6\sqrt{x-2}+x-2\]

OpenStudy (anonymous):

so one radical is gone, but one still remains now i combine like terms on the left, then isolate the radical so you can square again

OpenStudy (anonymous):

\[4x+1=7+x+6\sqrt{x-2}\] \[3x-6=6\sqrt{x-2}\]

OpenStudy (anonymous):

now you have to square one more time to get rid of the radical on the right don't forget to square the six too you get \[(3x-6)^2=36(x-2)\] and now you have a quadratic equation to solve

OpenStudy (anonymous):

4x+1=7+x+6√x-2 how did you get the 7? i thought it was 6

OpenStudy (anonymous):

fortunately this one is not too hard to solve \[9x^2-36x+36=36x-72\] \[9x^2-72x+108=0\] divide by 9

OpenStudy (anonymous):

did you understand how i got this line? \[4x+1=9+6\sqrt{x-2}+x-2\]

OpenStudy (anonymous):

the \(7\) came from the \(9-2\)

OpenStudy (anonymous):

ohh!! okay

OpenStudy (anonymous):

sorry i was a little lost on that one line, yeah i understand (:

OpenStudy (anonymous):

so you'd get x^2-8x+12?

OpenStudy (anonymous):

yes, now factor and solve and check your answers